OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote at y = 1 and a vertical asymptote at x = 4.

OpenStudy (anonymous):

what a great question what must the denominator look like if the vertical asymptote is \(x=4\) ?

OpenStudy (anonymous):

(x-4)?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

and horizonal 1/x?

OpenStudy (anonymous):

well, it could also be \(2x-8\) but why not keep it simple and write what you did

OpenStudy (anonymous):

so you know it looks like \[\frac{something}{x-4}\] all you need is the "something"

OpenStudy (anonymous):

the horizontal asymptote is \(y=1\) and the degree of the denominator is \(1\) so what must the degree of the numerator be?

OpenStudy (anonymous):

x-1?

OpenStudy (anonymous):

that would work, yes, but maybe not for the reason you think

OpenStudy (anonymous):

the degree of the numerator must be the same as the degree of the denominator, and the degree of the denominator is 1, so it could be \[\frac{x}{x-4}\]

OpenStudy (anonymous):

or it could be \[\frac{x-1}{x-4}\]if you like

OpenStudy (anonymous):

and that would answer both questions?

OpenStudy (anonymous):

either way the degree of the numerator must be 1, and the leading coefficient must also be 1 so you can use any example you like, so long as the denominator is \(x-4\) and the numerator starts with \(x\)

OpenStudy (anonymous):

your example would work fine so would \[\frac{x+5}{x-4}\] anything reallyl

OpenStudy (anonymous):

alright thanks again!